# MOWA: Multiple-in-One Image Warping Model

### v1

@@ Abstract
While recent image warping approaches achieved remarkable success on existing benchmarks, they still require training separate models for each specific task and cannot generalize well to different camera models or customized manipulations. To address diverse types of warping in practice, we propose a Multiple-in-One image WArping model (named MOWA) in this work. Specifically, we mitigate the difficulty of multi-task learning by disentangling the motion estimation at both the region level and pixel level. To further enable dynamic task-aware image warping, we introduce a lightweight point-based classifier that predicts the task type, serving as prompts to modulate the feature maps for better estimation. To our knowledge, this is the first work that solves multiple practical warping tasks in one single model. Extensive experiments demonstrate that our MOWA, which is trained on six tasks for multiple-in-one image warping, outperforms state-of-the-art task-specific models across most tasks. Moreover, MOWA also exhibits promising potential to generalize into unseen scenes, as evidenced by cross-domain and zero-shot evaluations. The code will be made publicly available.
Keywords:
Image Warping Multiple-in-One Model Prompt Learning
@@
1 Introduction
Image warping is essential in the field of computational imaging and computer vision, serving as the foundation for numerous applications, including image rectification [7, 31], image rectangling [12, 23], camera calibration [13, 1, 17], and 3D reconstructions [10], etc. By enabling the manipulation of image data through processes such as scaling, rotation, and sheering, they allow for the seamless integration of diverse visual elements and the correction of optical imperfections. Moreover, image warping is indispensable in developing augmented reality (AR) and virtual reality (VR) applications, where it helps create immersive and realistic environments by accurately mapping textures and images onto 3D models.
@@ 1 Introduction
In this work, we propose a Multiple-in-One image WArping model (named MOWA) [^1] We believe “multiple-in-one” is more rigorous than “all-in-one” or “universal” in our settings. to address various tasks in practice, as shown in Fig. 1. Specifically, we consider six representative types in the field of computational photography, namely stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks. Given the fact that learning different structures of motion is non-trivial in one model and motion representations differ significantly across various tasks, we propose to disentangle the motion estimation at both the region level and pixel level. In this hierarchical architecture, we first estimate the control points of the thin-plate spline (TPS) model [3] with increasing refined numbers, in which the feature maps are progressively warped and rectified. Such a representation excels in approximating complex motions at a region level and enables high flexibility to various motion structures. Subsequently, the warped feature maps are fed into the decoder to predict the residual pixel-level displacement, which further improves the warping results for each task, especially in the image boundaries and details.
To enable MOWA to explicitly discriminate diverse input types, a lightweight point-based classifier is devised. Adding an extra classification network based on the image features is a straightforward solution but brings high computation and storage costs. Noticing the motion structures in different warping tasks possess their specific distribution, we leverage the middle product of the image warping framework, i.e., region-level control points, to directly learn the task type. It achieves comparable performance while allowing significant parameter reduction compared to the image-based classifier since only a few 2D points are needed. Then, the soft task label predicted by this point-based classifier is used to modulate the feature maps in the decoder, dynamically boosting task-aware image warpings using a prompt learning module. Prompts are a set of learnable parameters that encapsulate essential discriminative information about different types of input, which empower a single model to efficiently traverse and harness its vast parameter space to accommodate various warping requirements.
In the experiments, we trained MOWA on six typical tasks for multiple-in-one image warping and demonstrated it outperforms state-of-the-art (SotA) task-specific models in most tasks. In addition, MOWA allows the ability to generalize to unseen scenes, as evidenced by cross-domain evaluation and zero-shot evaluation, indicating its robustness and adaptability across various scenarios. Our contributions can be summarized as follows:
@@
  • We propose to mitigate the difficulty of multi-task learning by decoupling the motion estimation in both the region level and pixel level.
  • A prompt learning module, guided by a lightweight point-based classifier, is designed to facilitate task-aware image warpings.
2 Related Work
@@
Excluding the customized manipulations, some special camera models can introduce geometric distortion onto the captured images, $e.g.$, radial distortion, rolling shutter distortion, and perspective distortion. The images’ semantic features significantly disobey the real-world rules due to those distortions. To address this issue, there is an exploration of distortion correction approaches [16, 2, 34, 7, 33, 21, 27, 30, 29, 40] aimed at warping the distorted input to a geometrically reasonable one. Particularly, regression-based methods [2, 27] learn the camera and distortion parameters from the input image and correct the distortion by simulating the imaging process of a predefined camera model. In contrast, reconstruction-based methods [16, 34, 7, 33, 30] directly learn the pixel-wise displacement between the distorted image and its ground truth, facilitating the model-free correction and enabling the end-to-end training.
@@
The above works achieve remarkable progress on various tasks, of which well-designed network architectures and tailored motion representations are studied. However, they need to train an individual model for each specific warping type and require the prior knowledge of the camera model or customized manipulations. In this work, we propose a multiple-in-one framework to involve these typical and practical image warping tasks.
3 Multiple-in-One Warping Model
3.1 Problem Definition
In this study, we consider six representative and practical image types in the field of computational photography, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos. These types are further classified into two groups. The first four types (stitched images, rectified wide-angle images, unrolling shutter images, and rotated images) struggle with irregular boundaries as the original images are manipulated by some customized operations, such as image stitching, distortion correction, and rotation. Therefore, image rectangling is proposed to reshape these irregular boundaries while keeping the distribution of content unchanged. The last two types (fisheye images and portrait photos) show inherently geometric distortions imaged by special camera models, such as radial distortion in fisheye images and perspective distortion in portrait photos. Correcting these distortions is crucial to scene understanding and aesthetic appreciation.
Figure 2: Overview of the proposed multiple-in-one image warping model (MOWA).
@@ 3.1 Problem Definition
Figure 2 shows the overall framework of the proposed MOWA. It takes the image and mask as input and estimates the TPS control points with increasingly refined numbers. In this region-level motion estimation, the feature maps are progressively warped and rectified. Subsequently, the warped features are fed into the decoder to predict a residual pixel-level motion. To enable task-aware and expandable capabilities, a lightweight point-based classifier and prompt learning module are designed. We elaborate on the details of each module as follows.
3.2 Motion Estimation Module
@@
The TPS transformation [3] stands out for its remarkable ability to model complex motions [6, 28, 15]. It is adept at performing image warping based on two sets of region-level control points, namely $\bm{Q}=[\bm{q}_{1},\bm{q}_{2},\cdots,\bm{q}_{N}]^{T}\in\mathbb{R}^{N\times 2}$ for the source image and $\bm{Q}^{{}^{\prime}}=[\bm{q}^{\prime}_{1},\bm{q}^{\prime}_{2},\cdots,\bm{q}^{\prime}_{N}]\in\mathbb{R}^{N\times 2}$ for the target image. To minimize the distortion of the source and target images, an energy term is introduced to penalize the Euclidean distance between the transformed source points $\mathcal{T}(\bm{q}_{i})$ and the target points $\bm{q}^{\prime}_{i}$, i.e., $\sum_{i=1}^{N}\lVert\mathcal{T}({\bm{q}_{i}})-{\bm{q}^{\prime}_{i}}\rVert_{2}^{2}$. This penalty results in a spatial deformation function parameterized by the control points, effectively capturing the intricate deformations across the image and maintaining the overall structural integrity. Specifically, the derived spatial deformation function can be expressed as follows:
@@
While TPS transformation is flexible to various tasks, it suffers from a limited degree of freedom to describe the detailed motions. To alleviate this limitation, we further complement the region-level motion representation with a pixel-level residual flow. Specifically, we first rectify the feature map $\mathcal{F}^{(T)}$ using the corresponding control points $\bm{q}^{(T)}$ in the last transformation head, and then feed the rectified feature map into a decoder network $f(\cdot)$ to predict the desirable residual flow. Like common U-Net architectures, the shallow features in the encoder are transited into the decoder using skip-connection. To eliminate the blur effect by multiple warpings (interpolation operation involved), we densify the TPS control points to pixel level and couple it with the residual flow to directly warp the input image. The final warping result can be formulated as follows:
$I^{\prime}=W(f[\mathcal{R}(\mathcal{F}^{(T)},\bm{q}^{(T)})]+\text{DE}[\bm{q}^{(T)}],I),$ (3)
@@
where $W(\cdot,\cdot)$ denotes the warping operation given the flow map and input image, $\text{DE}[\cdot]$ densifies the sparse control points to a dense flow map, which can be regarded as a special case of TPS upsampling layer $\text{UP}[\cdot]$.
3.3 Point-based Task Classifier
@@
It is straightforward to design a task classifier in terms of the input image. However, such a design brings an unavoidable issue of high computation complexity due to the redundant image features. Instead, we propose a lightweight task classifier based on the TPS points predicted by the motion estimation module. Our motivation stems from the fact that the motion structures in different tasks possess their specific distribution, which potentially exists in a 2D point space as shown in Fig. 3. To this end, we design a PointNet-like network [26] to predict the task type. Specifically, it takes the local coordinates of motion with the global image features (after maxpooling) by point-wise concatenation along the last dimension as input and outputs the soft task label, as shown in Fig. 2.
@@
In addition to the task classification function, our point-based task classifier can further improve image warping performance. This improvement is due to the high-level guidance provided by the task classifier to the motion estimation module through gradient back-propagation. Figure 3 (right) depicts a typical example, in which the fisheye rectification result shows a less distorted shape, and the predicted control points of the stitched image are more tightly aligned to the image boundary. Experiments demonstrate our point-based task classifier achieves comparable results while having less than ×$\times 50$ parameters compared with the image-based classifier. More quantitative results are presented in Sec. 4.3.
3.4 Prompt Learning Module
@@
Once the inputs are classified by the proposed point-based network, we leverage the predicted soft task label to modulate the feature maps in our MOWA. In particular, a prompt learning block is inserted into each layer in the decoder as a plug-and-play module. Prompt learning aims to tackle the challenge of generalizing in various image warping tasks by aiding the network in comprehending the specific task at hand. The prompts serve as a flexible and lightweight component to encode motion context across multiple scales within the image warping network.
@@
Assuming the task number is $N$, we introduce a set of learnable parameters as our prompts, namely $\{\bm{P}_{i}\}_{i=1}^{N}$. By denoting the predicted soft label by the task classifier as $\bm{\Phi}$, we modulate the feature maps $\mathcal{F}$ in the decoder network by the prompts as follows:
@@
where $\oplus$ represents the concat operation, $\text{Conv}_{1\times 1}$ is a convolution layer with $1\times 1$ kernel size aiming to reduce the channel dimension of concatenated features. By integrating prompts with the feature of MOWA, we can effectively enrich the representations with task-specific knowledge. This adaptive process empowers the network to efficiently warp images by dynamically influencing its behavior.
3.5 Training Loss
After predicting the TPS control points and the residual flow, the warped image can be obtained by Eq. (3). Following previous works [23, 18], we first exploit three losses to train our multiple-in-one image warping framework, e.g., image reconstruction loss $\mathcal{L}_{Rec}$, perceptual loss $\mathcal{L}_{Per}$, and inter-grid loss $\mathcal{L}_{Grid}$. The reconstruction loss and perceptual loss supervise the warped image at pixel level and feature level, respectively, while the inter-grid loss constrains the edges of two consecutive deformed grids to be co-linear. Considering the ground truth of warping flow is available in the dataset of portrait photos, we also add the reconstruction loss $\mathcal{L}_{Flow}$ on the predicted flow of the portrait correction task. Moreover, we provide middle-level supervision on the warped results from the TPS prediction heads with a set of exponentially growing weights. To train the point-based task classifier, the standard cross-entropy loss $\mathcal{L}_{Cls}$ is applied. Overall, the final loss can be expressed by:

### v2

@@ Abstract
While recent image warping approaches achieved remarkable success on existing benchmarks, they still require training separate models for each specific task and cannot generalize well to different camera models or customized manipulations. To address diverse types of warping in practice, we propose a Multiple-in-One image WArping model (named MOWA) in this work. Specifically, we mitigate the difficulty of multi-task learning by disentangling the motion estimation at both the region level and pixel level. To further enable dynamic task-aware image warping, we introduce a lightweight point-based classifier that predicts the task type, serving as prompts to modulate the feature maps for more accurate estimation. To our knowledge, this is the first work that solves multiple practical warping tasks in one single model. Extensive experiments demonstrate that our MOWA, which is trained on six tasks for multiple-in-one image warping, outperforms state-of-the-art task-specific models across most tasks. Moreover, MOWA also exhibits promising potential to generalize into unseen scenes, as evidenced by cross-domain and zero-shot evaluations. The code and more visual results can be found on the project page: https://kangliao929.github.io/projects/mowa/.
Index Terms:
Image Warping, Multiple-in-One Model, Prompt Learning.
@@
I Introduction
Image warping is essential in the field of computational imaging and computer vision, serving as the foundation for numerous applications, including image rectification [1, 2, 3, 4], image rectangling [5, 6, 7, 8], camera calibration [9, 10, 11, 12, 13], and 3D reconstructions [14, 15, 16, 17], etc. Enabling the manipulation of image data through processes such as scaling, rotation, and sheering allows for the seamless integration of diverse visual elements and the correction of optical imperfections. Moreover, image warping is indispensable in developing augmented reality (AR) and virtual reality (VR) applications [18, 19, 20], where it helps create immersive and realistic environments by accurately mapping textures and images onto 3D models.
@@ I Introduction
In this work, we propose a Multiple-in-One image WArping model (named MOWA) to address various tasks in practice, as shown in Fig. 1. Specifically, we consider six representative types in the field of computational photography, namely stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks.
Given the fact that learning different structures of motion is non-trivial in one model and motion representations differ significantly across various tasks, we propose to disentangle the motion estimation at both the region level and pixel level. In this hierarchical architecture, we first estimate the control points of the thin-plate spline (TPS) model [27] with increasing refined numbers, in which the feature maps are progressively warped and rectified. Such a representation excels in approximating complex motions at a region level and enables high flexibility to various motion structures. Subsequently, the warped feature maps are fed into the decoder to predict the residual pixel-level displacement, which further improves the warping results for each task, especially in the image boundaries and details.
To enable MOWA to explicitly discriminate diverse input types, a lightweight point-based classifier is devised. Adding an extra classification network based on the image features is a straightforward solution but brings high computation and storage costs. Noticing the motion structures in different warping tasks possess their specific distribution, we leverage the middle product of the image warping framework, i.e., region-level control points, to directly learn the task type. It achieves comparable performance while allowing significant parameter reduction compared to the image-based classifier since only a few 2D points are needed. Then, the task label predicted by this point-based classifier is used to modulate the feature maps in the decoder, dynamically boosting task-aware image warpings using a prompt learning module. Prompts are a set of learnable parameters that encapsulate essential discriminative information about different types of input, which empower a single model to efficiently traverse and harness its vast parameter space to accommodate various warping requirements.
In the experiments, we trained MOWA on six typical tasks for multiple-in-one image warping. Experimental results demonstrated that it outperforms state-of-the-art (SotA) task-specific models in most tasks, even with comparable parameters of the network. In addition, MOWA allows the ability to generalize to unseen scenes, as evidenced by cross-domain evaluation (unfamiliar domains) and zero-shot evaluation (unseen tasks), indicating its robustness and adaptability across various scenarios. Our contributions can be summarized as follows:
@@
  • We propose to mitigate the difficulty of multi-task learning by decoupling the motion estimation in both the region level and pixel level. Moreover, a prompt learning module, guided by a lightweight point-based classifier, is designed to facilitate task-aware image warpings.
  • We show that through multi-task learning, our framework develops a robust generalized warping strategy that gains improved performance across various tasks and even generalizes to unseen tasks.
The remainder of the paper is organized as follows: Section II reviews the related works. We then present the proposed multiple-in-one image warping framework in Section III. The experiments are provided in Section IV. Section V concludes this paper.
II Related Work
@@
Excluding the customized manipulations, some special camera models can introduce geometric distortion onto the captured images, $e.g.$, radial distortion, rolling shutter distortion, and perspective distortion. The images’ semantic features significantly disobey the real-world rules due to those distortions. To address this issue, there is an exploration of distortion correction approaches [31, 32, 37, 3, 38, 39, 40, 25, 41, 26, 4] aimed at warping the distorted input to a geometrically reasonable one. Particularly, regression-based methods [32, 40] learn the camera and distortion parameters from the input image and correct the distortion by simulating the imaging process of a predefined camera model. In contrast, reconstruction-based methods [31, 37, 3, 38, 25, 42] directly learn the pixel-wise displacement between the distorted image and its ground truth, facilitating the model-free correction and enabling the end-to-end training.
@@
The above works achieve remarkable progress on various tasks, of which well-designed network architectures and tailored motion representations are studied. However, they need to train an individual model for each specific warping type and require prior knowledge of the camera model or customized manipulations. In this work, we propose a multiple-in-one framework to involve these typical and practical image warping tasks. We address the challenge of learning different motion structures within a single model by employing a coarse-to-fine approach, progressively adding more TPS points to accurately fit the expected geometric distribution. To compensate additional degrees of freedom for TPS, our method further learns the residual flow based on the warped feature map, allowing for tuning of image boundaries and details in the final results.
@@
Figure 2: Overview of the proposed multiple-in-one image warping model (MOWA). It begins by taking an image and a mask as input to estimate the TPS control points with progressively refined precision. During such a region-level motion estimation, feature maps are incrementally warped and rectified. These warped features are then passed to the decoder to predict residual pixel-level motion. To ensure task awareness and expandability, a lightweight point-based classifier and a prompt learning module are designed. During inference, MOWA supports image warping for any resolution by scaling the predicted TPS control points and residual flow.
III Multiple-in-One Warping Model
III-A Problem Definition
In this study, we consider six representative and practical image types in the field of computational photography, including stitched images, rectified wide-angle images, unrolling shutter images, rotated images, fisheye images, and portrait photos, covering the mainstream practical image warping tasks. These types are further classified into two groups. The first four types (stitched images, rectified wide-angle images, unrolling shutter images, and rotated images) struggle with irregular boundaries as the original images are manipulated by some customized operations, such as image stitching, distortion correction, and rotation. Therefore, image rectangling is proposed to reshape these irregular boundaries while keeping the distribution of content unchanged. The last two types (fisheye images and portrait photos) show inherently geometric distortions imaged by special camera models, such as radial distortion in fisheye images and perspective distortion in portrait photos. Correcting these distortions is crucial to scene understanding and aesthetic appreciation.
Figure 2 shows the overall framework of the proposed MOWA. It takes the image and mask as input and estimates the TPS control points with increasingly refined numbers. In this region-level motion estimation, the feature maps are progressively warped and rectified. Subsequently, the warped features are fed into the decoder to predict a residual pixel-level motion. To enable task-aware and expandable capabilities, a lightweight point-based classifier and prompt learning module are designed. We elaborate on the details of each module in the proposed multiple-in-one image warping framework as follows.
III-B Motion Estimation Module
@@
Region-Level Motion Estimation. The TPS transformation [27] stands out for its remarkable ability to model complex motions [43, 44, 45]. It is adept at performing image warping based on two sets of region-level control points, namely $\bm{Q}=[\bm{q}_{1},\bm{q}_{2},\cdots,\bm{q}_{N}]^{T}\in\mathbb{R}^{N\times 2}$ for the source image and $\bm{Q}^{{}^{\prime}}=[\bm{q}^{\prime}_{1},\bm{q}^{\prime}_{2},\cdots,\bm{q}^{\prime}_{N}]\in\mathbb{R}^{N\times 2}$ for the target image. To minimize the distortion of the source and target images, an energy term is introduced to penalize the Euclidean distance between the transformed source points $\mathcal{T}(\bm{q}_{i})$ and the target points $\bm{q}^{\prime}_{i}$, i.e., $\sum_{i=1}^{N}\lVert\mathcal{T}({\bm{q}_{i}})-{\bm{q}^{\prime}_{i}}\rVert_{2}^{2}$. This penalty results in a spatial deformation function parameterized by the control points, effectively capturing the intricate deformations across the image and maintaining the overall structural integrity. Specifically, the derived spatial deformation function can be expressed as follows:
@@
Pixel-Level Motion Estimation. While TPS transformation is flexible and adaptable to various tasks, it is limited in its ability to describe detailed motions due to its restricted degree of freedom. To alleviate this limitation, we further complement the region-level motion representation with a pixel-level residual flow. Specifically, we first rectify the feature map $\mathcal{F}^{(T)}$ using the corresponding control points $\bm{q}^{(T)}$ in the last transformation head, and then feed the rectified feature map into a decoder network $\mathcal{D}[\cdot]$ to predict the desirable residual flow. Like common U-Net architectures, the shallow features in the encoder are transited into the decoder using skip-connection. To eliminate the blur effect by multiple warpings (interpolation operation involved), we densify the TPS control points to pixel level and couple it with the residual flow to directly warp the input image $\mathcal{I}$. The final warping result $\mathcal{I}^{\prime}$ can be obtained by:
$\mathcal{I}^{\prime}=\mathcal{W}\left(\mathcal{D}[\mathcal{R}(\mathcal{F}^{(T)},\bm{q}^{(T)})]+\text{DE}[\bm{q}^{(T)}],\mathcal{I}\right),$ (3)
@@
where $\mathcal{W}(\cdot,\cdot)$ denotes the warping operation given the flow map and input image, $\text{DE}[\cdot]$ densifies the sparse control points to a dense flow map, which can be regarded as a special case of TPS upsampling layer $\text{UP}[\cdot]$. Unlike previous works tailored for specific warping tasks, our method unifies motion representation across various tasks at both the region and pixel levels. The experimental results are demonstrated in Section IV-C.
III-C Point-based Task Classifier
@@
When learning various image warping tasks simultaneously, a task classifier is crucial for efficiently routing inputs to their respective task-specific components, optimizing resource use, and enhancing model performance. It is straightforward to design a task classifier in terms of the input image. However, such a design brings an unavoidable issue of high computation complexity due to the redundant image features. Instead, we propose a lightweight task classifier based on the TPS points predicted by the motion estimation module. Our motivation stems from the fact that the motion structures in different tasks possess their specific distribution, which potentially exists in a point space as shown in Fig. 3. To this end, we design a PointNet-like network [46, 47] to predict the task type. Specifically, as shown in Fig. 2, it takes the local coordinates of motion with the global image features (after maxpooling) by point-wise concatenation along the last dimension as input and outputs the soft task label $\bm{\Phi}$. We can formulate this point-based task classifier as follows:
$\mathcal{F}_{g}=f\left[\text{MaxPool}(\mathcal{F})\right],$ (4)
$\bm{\Phi}=\text{Softmax}\left(f^{\prime}[\bm{q}^{(T)}\oplus\text{R}(\mathcal{F}_{g},[1,H\times W])]\right),$ (5)
where $f$ and $f^{\prime}$ are the fully connected layers to decrease the dimensions of features and learn the abstract concepts. R denotes replicating $\mathcal{F}_{g}$ to the same shape of the predicted motion coordinates $\bm{q}^{(T)}$ and $\oplus$ is the point-wise concatenation. Experiments demonstrate our point-based task classifier achieves comparable results while having less than ×$\times 50$ parameters compared with the image-based classifier.
@@
In addition to the task classification function, our point-based task classifier can further improve image warping performance. This improvement is due to the high-level guidance provided by the task classifier to the motion estimation module through gradient back-propagation. Figure 3 (right) depicts a typical example, in which the fisheye rectification result shows a less distorted shape, and the predicted control points of the stitched image are more tightly aligned to the image boundary. Compared to the vanilla baseline, the proposed point-based task classifier achieves an average improvement of $+0.35$ in PSNR metrics across various image warping tasks. More quantitative results are presented in Section IV-C.
III-D Prompt Learning Module
@@
Once the inputs are classified by the proposed point-based network, we leverage the predicted task label to modulate the feature maps in the network. In particular, a prompt learning block is inserted into each layer in the decoder as a plug-and-play module. Prompt learning aims to tackle the challenge of generalizing in various image warping tasks by aiding the network in comprehending the specific task at hand. The prompts serve as a flexible and lightweight component to encode motion context across multiple scales within the image warping network.
@@
Assuming the task number is $N$, we introduce a set of learnable parameters as our prompts, namely $\{\bm{P}_{i}\}_{i=1}^{N}$. By denoting the predicted task label by the task classifier as $\bm{\Phi}$, we modulate the feature maps $\mathcal{F}$ in the decoder network by the prompts as follows:
@@
where $\oplus$ represents the concat operation, $\text{Conv}_{1\times 1}$ is a convolution layer with $1\times 1$ kernel size aiming to reduce the channel dimension of concatenated features.
By integrating the learnable prompts with the features of the warping model, we can significantly enrich the representations with task-specific knowledge. Unlike pre-defined and fixed prompts, our adaptive approach enables the network to dynamically influence its behavior, resulting in more efficient and precise image warping. This adaptive process not only enhances the flexibility of the model but also improves its ability to generalize across different tasks and datasets. More analysis on the multi-task learning and effectiveness of the proposed prompt learning are demonstrated in Section IV-D.
III-E Training Loss
After predicting the TPS control points and the residual flow, the warped image can be obtained by Eq. (3). Following previous works [6, 21], we first exploit three losses to train our multiple-in-one image warping framework, e.g., image reconstruction loss $\mathcal{L}_{Rec}$, perceptual loss $\mathcal{L}_{Per}$, and inter-grid loss $\mathcal{L}_{Grid}$. The reconstruction loss and perceptual loss supervise the warped image at the pixel level and feature level, respectively. The inter-grid loss constrains the edges of two consecutive deformed grids $\{\vec{e}_{t1},\vec{e}_{t2}\}$ to be co-linear:
$\mathcal{L}_{Grid}=\frac{1}{M}\sum_{\{\vec{e}_{t1},\vec{e}_{t2}\}\in m}(1-\frac{\langle\vec{e}_{t1},\vec{e}_{t2}\rangle}{\parallel\vec{e}_{t1}\parallel\cdot\parallel\vec{e}_{t2}\parallel}).$ (7)
Here, $M$ represents the number of tuples of two successive edges in a mesh $m$. When maximizing the above cosine representation, the corresponding two edges become collinear. Consequently, the loss reaches its minimum, ensuring the image content remains consistent.
Considering the ground truth of warping flow is available in the training dataset of portrait photos, we also add the reconstruction loss $\mathcal{L}_{Flow}$ on the predicted flow of the portrait correction task. Moreover, we provide middle-level supervision on the warped results from the TPS prediction heads with a set of exponentially growing weights. To train the point-based task classifier, the standard cross-entropy loss $\mathcal{L}_{Cls}$ is applied. Overall, the final loss can be expressed by: